# Improving User Mental Models of XAI Systems with an Inclusive Design Approach

### v2

@@ 3.1. Experimental Design
We set up a between-subject study in a lab to evaluate the two versions of Team Game’s XAI prototype.
@@ 3.1.1. Participants
To recruit participants, we distributed flyers through our university’s internal mailing lists and posted them around campus. The recruitment emails and flyers had the invitation link to a Qualtrics survey, where we collected demographic information and asked two eligibility questions to make sure our participants were at least 18 years of age and had not taken any AI courses as we wanted to recruit people with no AI background. Based on these criteria, we dropped 41 responses and sent out emails to the remaining 110 eligible respondents individually. Ultimately, 77 respondents accepted the invitations and 69 participated in our study. These 69 participants came from diverse backgrounds, spanning different genders, age groups, and college statuses (all summarized in Table 2), as well as 39 unique college majors.
@@ 3.1.2. Procedures
We first assigned participants to sessions based on their availability and then randomly assigned each session to one of two versions of Team Game’s game prototype: the “Original” and the “Post-GenderMag”. The Original sessions included 34 participants, while the Post-GenderMag sessions had 35 participants. Each session lasted 2 hours and had 2 to 6 participants. We conducted a total of 17 sessions from August 18 to August 28, 2023.
Upon arriving at the lab, participants filled out the GenderMag survey (provided in the supplemental document of (Vorvoreanu et al., 2019)) in which participants answered Likert-scale items about their problem-solving styles. Then we began the session with a 25-minute hands-on tutorial covering the game, agents, explanations, and participants’ tasks. Afterward, participants went through the first three moves of a practice game to get familiar with the prototype and the study questions.
In the main part of the study session, participants observed 3 games of Tic-Tac-Toe played by 2 AI agents: Agent Blue (X) and Agent Red (O). Before Agent Blue made a move, participants predicted the move, provided reasoning, and their confidence in it. After the move, they identified flaws and suggested changes. After each game, they answered two After-Action Review for AI (AAR/AI) questions. Once all three games concluded, they completed post-task questions covering agent descriptions, explanation feedback, and subjective workload assessment using NASA TLX (NASA, [n. d.]). Table 3 lists all the questions.
@@
- Explanation feedback  [^1] Participants provided feedback for each of the three explanations, BTW, STT, and OTB. They responded on a 7-point Likert-scale from Very Strongly Disagree to Strongly Agree.: i) the explanation helped in understanding Agent Blue, and ii) the explanation was sufficiently detailed.
@@
From these questions, we have not used the responses from “Please describe the steps in the agent’s approach or method…”. When the user study was over, we found that Agent Blue’s noise levels were different between each game. This means participants would be describing multiple agents rather than one. Despite Team Game’s lead (who created the agent) suggesting that the noise differences didn’t affect the agent’s behavior much, we excluded these responses from our analysis to ensure clarity in our evaluation of users’ understanding of AI.
3.1.3. Qualitative Analysis Methods
After all the study sessions were over, we used affinity diagramming to organize the participant responses into concepts participants expressed. Five of the researchers independently marked participants’ responses that described some aspect of their mental model of the AI agents. We then discussed the results as a group and developed the initial concept codeset. We also met with Team Game’s lead (the subject matter expert who had created the AI agents) to further modify the initial concept codeset. Finally, the concept codeset had 35 concept codes divided into four groups: “Prediction Why”, “Strategy Should Be”, “Flaws”, and “HowAgentBehaves” (Table LABEL:tab:codes_definitions). Four of the researchers (R1, R2, R3, and R4) then worked in pairs of two as per Table 4 to independently code ¿=20% of each of the four concept code groups. Table 4 lists the inter-rater reliability rates (IRR) (Stemler, 2004) of the agreement tests. Given the high IRRs, one researcher from each coder pair finished the rest of the coding for their assigned concept code group.
Table 4. Interrater-reliability results (IRR) for each concept code group
@@ 3.2. Where the Two Versions of the Prototype Came From
Before our study, Team Game used GenderMag to find and fix inclusivity issues in their game prototype. During the “find” sessions, Team Game identified inclusivity issues within the Original prototype (Figure 1). In the “fix” sessions, they proposed design solutions to potentially address these issues. These sessions took place over a period of five weeks via Zoom.
Figure 1. The Original prototype during a game. Columns are labeled A-I and rows are labeled 1-4. Squares can be identified using their column name followed by the row number (e.g., E1). The game progression is controlled by clicking the prior move and next move buttons below the gameboard.
Team Game started the find sessions by customizing the Abi persona, which they then used during the GenderMag walkthrough to identify inclusivity issues within their prototype. They identified issues that spanned all of GenderMag’s five problem-solving styles outlined in Table 1. Team Game engaged in find sessions totaling about five hours.
In the fix sessions, Team Game systematically reviewed all identified inclusivity issues from the find sessions spending about eight hours in total. Team Game devised 15 proposed fixes during these sessions. Table 5 presents two of these fixes which we later utilized to illustrate our study findings. All the 15 fixes are listed in Table LABEL:tab:fixes_all in the appendix.
Table 5. Two fixes, why Team Game fixed them (problem-solving styles impacted), and the changes made for the fixes
@@
Rest of the Fixes are in Table LABEL:tab:fixes_all.
We then implemented 13 of Team Game’s 15 fixes. We could not implement Fix-14 and Fix-15 due to their technical difficulties and our time constraints. Figure 2 shows the Post-GenderMag version of Team Game’s prototype.
Figure 2. The Post-GenderMag prototype during a game. Some fixes resulted in new additions such as the (A) Game History, (B) Game Log, and (C) Top 5 Moves. These are a few of the features that differentiate the Post-GenderMag version at this state of the game from the Original version (Figure 1) at the same state.
We call the original version “Original” and the redesigned version “Post-GenderMag” as its changes came from the GenderMag-driven cognitive walkthroughs. With these two versions, we aimed to address our research questions about Team Game’s GenderMag-driven fixes’ effects on participants’ mental models, prediction performance, and explanation ratings.
@@ 3.3. Game Overview and Prototype
Team Game was working in the sequential decision domain of MNK games, based on Dodge et al.’s original prototype (Dodge et al., 2022). MNK games are an abstraction of Tic-Tac-Toe where M and N represent the dimensions of the gameboard ($M\times N$) and the goal is to get a sequence of length K. To simplify the prototype presented in Dodge et al. (Dodge et al., 2022) and enable Team Game to make changes, we re-created the original prototype in a web-based format using Svelte (JavaScript framework) seen in Figure 1. In this MNK game, the gameboard—on the left half of the prototype—has 9 columns and 4 rows (36 squares) where the winning condition is getting 4 in-a-row (9-4-4). The game is played by two AI agents, Agent Blue (X) and Agent Red (O), that use a convolutional neural network to predict outcome tuples $O=(Win\%,Loss\%,Draw\%)$ for each square on the gameboard before making a move. Each square is then assigned a “Score” using the function $Win\%-Loss\%$. The agents make moves based on the scores. To the right of the gameboard, the prototype provides three explanations for Agent Blue’s move but not for Agent Red. This prototype had the same explanations presented in Dodge et al. (Dodge et al., 2022), namely Scores Best-To-Worst (BTW), Scores Through Time (STT), and Scores On-the-Board (OTB).
3.3.1. Explanation Type: Scores Best-to-Worst (BTW)
@@
The Scores Best-to-Worst (BTW) explanation presents all of Agent Blue’s predicted scores for the 36 gameboard squares ranked from the best move to the worst move on a graph (Figure 3). The X-axis is the order of scores, from the best move on the left to the worst moves on the right and the Y-axis represents the score from -100% to 100%. Agent Blue assigns very low scores for the illegal moves (moves that are already taken). After each move by Agent Blue, a new series of rectangles will appear in the explanation and the prior set of scores will get a lighter color. The most recent data series will have the darkest color and can be interacted with by hovering. Hovering over a score in the most recent data series will highlight its corresponding square on the gameboard and its past scores in all the explanations. Hovering over a square on the gameboard will highlight its corresponding scores in BTW (and also in other 2 explanations) over the past moves.
@@
Figure 3. BTW in the Original prototype. Three data series of scores by Agent Blue after the agent made its third move. The blue data series is the most recent scores the agent calculated.
3.3.2. Explanation Type: Scores Through Time (STT)
@@
The Scores Through Time (STT) explanation shows how the distribution of scores changes throughout the game whenever Agent Blue makes a move (Figure 4). The X-axis is time (decision points) and the Y-axis is the score. When Agent Blue makes a move, a new column of scores appears. The score with the maximum score in the column is highlighted in a darker color for the current move. Hovering over a score in the column highlights its corresponding square in the gameboard and the score of that square in prior decisions. Like the prior explanation BTW, only the most recent series of scores can be hovered over. Hovering over a gameboard square will highlight its score for every decision through time.
@@
Figure 4. STT in the Original prototype. Agent Blue has played 3 moves so there are 3 columns and the number under each column represents the move number.
3.3.3. Explanation Type: Scores On-the-Board (OTB)
The last explanation, Scores On-the-Board (OTB), shows the scores through time at the level of each square where each square has its own version of STT (Figure 5). Each little dot shows Agent Blue’s evaluation of that square over different moves or decision points. Squares that have been used for moves have a highlighted border in the same color (blue or red) as the agent that picked the move. Hovering over a square in the gameboard highlights the corresponding chart in the explanation and vice versa. A new data point is added to each square’s chart every time Agent Blue makes a move.
@@ 3.3.3. Explanation Type: Scores On-the-Board (OTB)
Figure 5. OTB in the Original prototype. The square highlighted in yellow is the most recent move by Agent Blue which has a high score.
3.4. Measuring Mental Models for this Game
3.4.1. Mental Model Rubric
@@
We needed a mental model rubric to calculate the mental model score of each participant. From the four concept code groups of the initial codeset, Team Game’s lead, who created the AI agents, selected two groups, “Flaws” and “HowAgentBehaves” for the mental model rubric (Table 6).
“Flaws”: The agent creator picked the following three concepts under “Flaws” for the rubric:
– Square X is miscalculated.
– Agent miss selection.
– Agent chose the wrong action.
@@
The agent creator pointed out that to identify flaws, participants needed to understand the scoring system, the agent’s decision-making process, and the gameboard. They also mentioned both local and global understanding of the Agents’ moves were necessary for detecting flaws. For these reasons, the agent creator picked the above three concepts that reflected Agent Blue’s decision-making “Flaws”. They assigned positive weights to these concepts in the mental model rubric.
“HowAgentBehaves”: The agent creator picked seven concepts from the “HowAgentBehaves” group for the mental model rubric and divided them into two subgroups. “Agent” mentioned in the following concepts is Agent Blue and “Opponent” is Agent Red:
i) HowAgentBehaves concepts which are Correct:
– Agent picks high score.
– Agent doesn’t pick high score.
– Agent doesn’t consider opponent’s moves.
ii) HowAgentBehaves concepts which are INcorrect:
– Opponent is random/illogical.
– Agent is random/illogical.
– Agent prolongs game.
– Agent is/can learn
@@
The agent creator considered that these seven concepts reflected on participants’ perceptions of the AI agents’ gameplay, behavior, priorities, strengths, and weaknesses. According to the agent creator, these concepts would cover the required understanding to evaluate a participant’s comprehensive mental model of the AI agents, justifying their integration into the mental model rubric. The agent creator also pointed out that mentioning the correct concepts from subgroup-i indicated participants’ better understanding of the AI agents, while INcorrect concepts from subgroup-ii indicated a lack of understanding. That’s why they assigned positive weights for subgroup-i and negative weights for subgroup-ii in the mental model rubric.
@@
Table 6. Concepts used in the mental model rubric, and the weights/weight ranges Team Game assigned
Concept & Definition Weight Range
Square “X” is miscalculated: Participant says score is wrong/agent miscalculated the score/there are prediction flaws and they specified which particular square they were referring to. [6, 10]
Agent miss-selection: Participant says agent selected the wrong square because it is not the high/highest score (not following its reasoning). 7
Agent chose the wrong action: Participant says agent missed a win opportunity, missed a critical move, or missed an opportunity to block. [11, 15]
Agent picks high score: Participant says agent is picking the square with the highest score or within the top 5. 2
Agent doesn’t pick high score: Participant says agent is not picking the square with the highest score or within the top 5. [6, 10]
Agent doesn’t consider opponent’s moves: Participant says agent does not consider the opponent or consider their effect. [11, 15]
Opponent is random/illogical: Participant says red is dumb, not smart, making weird moves, illogical, random. [-2, -1]
Agent is random/illogical: Participant says agent is generally random, behaving illogically, making weird moves, or doesn’t know what it is doing. [-5, -3]
Agent prolongs game: Participant says agent prolongs/stalls/lengthens the game (e.g., not doing finishing or winning move). [-5, -3]
Agent is/can learn: Participant says agent maybe/is learning. -15

### v3

@@ 3.1. Experimental Design
To investigate our research questions, we ran a between-subjects lab study to evaluate the Post-GenderMag vs. Original versions of Team Game’s XAI prototype.
@@ 3.1.1. Participants
To recruit participants, we distributed flyers through our university’s internal mailing lists and posted them around campus. The recruitment emails and flyers linked to a Qualtrics survey, which collected demographic information and asked two eligibility questions to make sure our participants were at least 18 years of age and had not taken any AI courses as we wanted participants with no AI background. Based on these criteria, we dropped 41 responses and sent out emails to the remaining 110 eligible respondents individually. Ultimately, 77 respondents accepted the invitations and 69 participated in our study. These 69 participants came from diverse backgrounds, spanning different genders, age groups, college statuses (all summarized in Table 2), and 39 unique college majors.
@@ 3.1.2. Procedures
We first assigned participants to sessions based on their availability and then randomly assigned each session to one of two versions of Team Game’s prototype: the “Original” and the “Post-GenderMag”. The Original sessions included 34 participants and the Post-GenderMag sessions had 35 participants. Each session lasted 2 hours and had 2 to 6 participants. We conducted a total of 17 sessions.
Upon arriving at the lab, participants filled out the GenderMag survey (Hamid et al., 2024) (using the version in Vorvoreanu et al. (Vorvoreanu et al., 2019)’s supplemental document), in which participants answered Likert-scale questions about their problem-solving styles. Then we began the session with a 25-minute hands-on tutorial covering the game, agents, explanations, and participants’ tasks. Participants then went through the first three moves of a practice game to get familiar with the prototype and the tasks.
For the main tasks, participants observed three games played by two AI agents: Agent Blue and Agent Red. Before Agent Blue made a move, participants predicted the move, provided reasoning, and rated their confidence level. After the move, they identified flaws and suggested changes. After each game, they answered two After-Action Review for AI (AAR/AI) questions (Dodge et al., 2021b). Once all three games concluded, they completed post-task questions covering agent descriptions, explanation feedback, and subjective workload assessment using NASA TLX (NASA, [n. d.]). Table 3 lists all the questions.
While designing the questionnaire, we used some of the mental model elicitation methods recommended by Hoffman et al. (Hoffman et al., 2019) such as “Prediction Task” for measuring participants’ ability to predict, “Task Reflection task” for participants’ reasoning, “Glitch Detection Task” for participants’ ability to find flaws, and “Self-Explanation Task” for participants’ understanding of AI.
@@
- Explanation feedback*: i) the explanation helped in understanding Agent Blue, and ii) the explanation was sufficiently detailed.
@@
*Participants provided feedback for each of the three explanations, BestToWorst, ScoresThroughTime, and OnTheBoard. They responded on a 7-point Likert-scale from Very Strongly Disagree to Very Strongly Agree used to compute ratings.
From Table 3’s questions, we discarded responses from “Please describe the steps in the agent’s approach or method…”, because Agent Blue’s noise levels differed. In games 1 and 3, the agent had the same noise levels, but different levels in game 2. This means participants would be describing different agents rather than one. To mitigate this threat, we excluded these responses from our analysis, despite Team Game’s lead (who created the agent) suggesting that the noise differences didn’t affect the agent’s behavior much.
3.2. Qualitative Analysis Methods
When all the study sessions were over, we used affinity diagramming to organize participant responses into concepts that the participants expressed. Five of the researchers independently marked participants’ responses that described some aspect of their mental model of the AI agents. We then discussed the results as a group and developed the initial concept codeset—further modifying it with input from Team Game’s lead (the subject matter expert who created the AI agents). Finally, the concept codeset had 35 concept codes divided into four groups: “Prediction Why”, “Strategy Should Be”, “Flaws”, and “How Agent Behaves” (Appendix’s Table LABEL:tab:codes_definitions). Four of the researchers (R1, R2, R3, and R4) then worked in pairs of two as per Table 4 to independently code ¿=20% of each of the four concept code groups. Table 4 lists the inter-rater reliability rates (IRR) (Stemler, 2004) of the agreement tests. Given the high IRRs, one researcher from each coder pair finished the rest of the coding for their assigned concept code group.
Table 4. Inter-rater reliability results (IRR) for each concept code group
@@ 3.2. Qualitative Analysis Methods
3.3. Statistical Analysis Methods
We used several statistical methods for comparing our participants. To compare participants’ mental model concepts scores and average explanation ratings we used Student’s one-tailed t-tests. One-tailed t-tests are suitable for detecting effects occurring only in one direction (Frost, 2021), whereas two-tailed tests are for both positive and negative effects. Considering these attributes of the two tests, we turned to the literature related to our study to investigate whether a one-tailed or two-tailed approach was more appropriate. For mental model concepts scores and explanation ratings, the literature points to one direction only, unanimously reporting GenderMag to improve the user experience of people with diverse problem-solving styles  (Guizani et al., 2022; Vorvoreanu et al., 2019; Burnett et al., 2017; Hilderbrand et al., 2020; Blackmon et al., 2003; Green et al., 2000). Thus, we used one-tailed t-tests when we hypothesized that inclusivity fixes would bring positive effects. For other comparisons, where we did not have reason to hypothesize that the effect can exist only in one direction, we used Student’s two-tailed t-tests. For the comparisons, we used Student’s one-tailed or two-tailed t-tests as appropriate when all three assumptions of independent samples t-test—1) independent observations, 2) normal distribution, and 3) equal variance were met. The independent observation condition was always true as each group was independent and the participants used the prototype individually with no interaction between them. When the normality distribution failed from a Shapiro-Wilk test, we conducted a non-parametric Wilcoxon Rank-Sum test to compare participants. Lastly, if equal variance failed from an F-test, we conducted a Welch’s t-test for the comparison.
We used mixed-design ANOVA for analyses that required investigating both within-subject and between subject-factors. To investigate if explanation usage could predict participants’ mental model concept scores, we used linear regression.
@@ 3.4. Where the Two Versions of the Prototype Came From
Before our study, Team Game had used GenderMag to find and fix inclusivity issues in their game prototype (Anderson, 2024). This was split into “find” and “fix” sessions on the prototype (Figure 2). During these sessions, Team Game evaluated their prototype from the perspective of GenderMag’s Abi persona as Abi represents the group of users who are most often left out of the technology development process. A snippet of what Team Game found and their proposed fixes is in Table 5. Table LABEL:tab:fixes_all in the Appendix has the complete list.
Figure 2. Original prototype (we implemented via Svelte) during a game. Participants controlled the game progression with the “Prior Move” and the “Next Move” buttons. Three explanations to the right provided details on Agent Blue (X)’s decisions.
Table 5. Two fixes, why Team Game fixed them (problem-solving styles impacted), and the changes made for the fixes. The rest of the fixes are in the Appendix’s Table LABEL:tab:fixes_all.
@@
Fix-2: Exact values and Exact actions “She is not clear on what she should do…..will not take this step.” PSS: Risk, Learn, Info, SE, Motiv. – Showed exact win%, loss%, and drawfor every rectangle in the explanations using tooltips. – Added tooltips that informed about the exact actions offered by the next and the prior move buttons.
@@
Figure 3. Post-GenderMag prototype during a game at the same state as the Original prototype as in Figure 2. Some fixes resulted in new additions such as the (A) Game History, (B) Game Log, and (C) Top 5 Moves. The Appendix’s Table LABEL:tab:fixes_all enumerates all differences between the Post-GenderMag version and the Original version.
We call the original version “Original” and the redesigned version “Post-GenderMag” (as its changes came from the GenderMag-driven walkthroughs). With these two versions, we aimed to address our research questions about GenderMag-driven inclusivity fixes’ impacts on participants’ mental model concepts scores, prediction accuracy, and inclusivity.
@@ 3.5. Game Overview and Prototype
Team Game was working in the sequential decision domain of MNK games, using a version of Dodge et al.’s prototype (Dodge et al., 2022). MNK games are an abstraction of Tic-Tac-Toe where M and N represent the dimensions of the gameboard ($M\times N$) and the goal is to get a sequence of length K. To enable Team Game to make changes in a lightweight way, we simulated the original prototype presented in Dodge et al. (Dodge et al., 2022) using a web-based platform using Svelte (JavaScript framework), shown in Figure 2. In this MNK game, the gameboard—on the left half of the prototype—has 9 columns and 4 rows (36 squares) where the winning condition is getting 4 in-a-row (9-4-4). Two AI agents, Agent Blue (X) and Agent Red (O) play the game and use a convolutional neural network to predict outcome tuples $O=(Win\%,Loss\%,Draw\%)$ for each square on the gameboard before making a move. The agents then assign a “Score” for each square using the function $Win\%-Loss\%$ and make moves based on the scores. To the right of the gameboard, the prototype provides three explanations for Agent Blue’s move but not for Agent Red. The prototype used the explanations presented in Dodge et al. (Dodge et al., 2022), namely Scores Best-To-Worst (BestToWorst), Scores Through Time (ScoresThroughTime), and Scores On-the-Board (OnTheBoard).
3.5.1. Explanation Type: Scores Best-to-Worst (BestToWorst)
@@
The Scores Best-to-Worst (BestToWorst) explanation presents all of Agent Blue’s predicted scores for the 36 gameboard squares ranked from the best move to the worst move on a graph (Figure 4). The X-axis orders scores in descending order with the best moves (higher score) on the left. The Y-axis shows the score value from -100% to 100%. Agent Blue assigns the lowest scores for illegal moves (i.e., moves that are already taken). With each move by Agent Blue, a new series of scores appears in a dark color in the graph, and the prior set of scores turns lighter in color. Hovering a mouse cursor over a score in the most recent data series highlights its corresponding past scores and gameboard square. Similarly, hovering over a gameboard square highlights its corresponding scores in the explanations.
@@
Figure 4. BestToWorst in the Original prototype. Agent Blue has made three moves so there are three series of scores. The most recent highest scoring square (blue series) is highlighted along with its past scores.
3.5.2. Explanation Type: Scores Through Time (ScoresThroughTime)
@@
The Scores Through Time (ScoresThroughTime) explanation shows the changes in score distribution over the course of the game (Figure 5: Left). The X-axis is time (decision points) and the Y-axis is the score. Each time Agent Blue makes a move, a new column appears in ScoresThroughTime with the agent’s predicted scores for all 36 gameboard squares. Each column organizes the scores from the highest at the top (marked in a darker color) to the lowest at the bottom. ScoresThroughTime has the same highlight interactions as BestToWorst.
@@
Figure 5. Left: ScoresThroughTime in the Original prototype. Agent Blue has played 3 moves so there are 3 columns of scores and the number under each column represents the move number.
@@
Right: OnTheBoard in the Original prototype. The square highlighted in yellow is the most recent move by Agent Blue which has a high score.
3.5.3. Explanation Type: Scores On-the-Board (OnTheBoard)
The last explanation, Scores On-the-Board (OnTheBoard), shows a mini-version of the ScoresThroughTime chart for each square on the board (Figure 5: Right). Each dot shows Agent Blue’s predicted score for that square over different moves. When Agent Blue makes a move, OnTheBoard adds a new data point to each square’s chart. If a square is occupied, it has a highlighted border in the same color (blue or red) as the occupying agent. Hovering over a square in the gameboard highlights the corresponding chart in the explanation and vice versa.
3.6. Measuring Mental Model Concepts for this Game
3.6.1. Mental Model Concepts Rubric
@@
We needed a mental model concepts rubric to calculate the mental model concepts score of each participant. From the four concept code groups of the codeset, Team Game’s lead, who had created the AI agents, selected from the list in the Appendix’s Table LABEL:tab:codes_definitions, “Flaws” and “How Agent Behaves” for the mental model concepts rubric (Table 6).
@@
The agent creator pointed out that to identify flaws, participants needed to understand the scoring system, the agent’s decision-making process, and the gameboard. They also mentioned both local and global understanding of the Agents’ moves were necessary for detecting flaws. For these reasons, the agent creator picked the three “Flaws” concepts (top three rows in Table 6) that reflected Agent Blue’s decision-making “Flaws” and assigned positive weights for these concepts in the mental model concepts rubric.
@@
The agent creator also selected the seven “How Agent Behaves” concepts (Table 6) to capture participants’ perceptions of the AI agents’ behaviors, including agents’ priorities, strengths, and weaknesses. According to the agent creator, these concepts would cover the required understanding to evaluate a participant’s comprehensive mental model of the AI agents, justifying the concepts’ integration into the mental model concepts rubric. The agent creator also pointed out that mentioning the correct concepts (top three rows in Table 6’s “How Agent Behaves” section) indicated participants’ better understanding of the AI agents. In contrast, the INcorrect concepts (bottom four rows) indicated a lack of understanding. Thus, the agent creator assigned positive weights for the correct concepts and negative weights for the incorrect ones in the mental model concepts rubric.
@@
Table 6. Concepts used in the mental model concepts rubric and their weights/weight ranges. $+$ sign indicates correct concepts with positive weights and -- sign indicates incorrect concepts with negative weights.